### v2

@@ 1.1 Literature Review
Learning for multi agent control problems is a practically relevant and a challenging problem where there has been as a growing interest in recent years. A general solution methodology for multi-agent control problems is difficult to obtain and the solution, in general, is intractable except for special information structures between the agents. We refer the reader to the survey paper by [43] for a substantive summary of learning methods in the context of multi-agent decision making problems.
@@
In this paper, we study a particular case of multi-agent problems in which both the agents and their interactions are symmetric and homogeneous. For these mean-field type decision making problems, the agents are coupled only through the so-called mean-field term. These problems can be broadly divided into two categories; mean-field game problems where the agents are competitive and interested in optimizing their self objective functions, and mean-field control problems, where the agents are interested in a common objective function optimization. We cite some papers by [21, 12, 11, 42, 26, 2, 18, 19, 24, 35, 40, 38, 39] and references therein, for papers in mean-field game setting. We do not discuss these in detail as our focus will be on mean-field control problems which are significantly different in both analysis and the nature of the problems of interest.
@@
For mean-field control problems, where the agents are cooperative and work together to minimize (or maximize) a common cost (or reward) function, see [8, 17, 30, 14, 36, 13, 20, 7, 10] and references therein for the study of dynamic programming principle and learning methods in continuous time. In particular, we point out the papers [28, 17] which provide the justification for studying the centralized limit problem by rigorously connecting the large population decentralized setting and the infinite population limit problem.
@@
For papers studying mean-field control in discrete time, we refer the reader to [33, 5, 22, 23, 32, 15]. [33, 32] study existence of solutions to the control problem in both infinite and finite population settings, and they rigorously establish the connection between the finite and infinite population problems. [5] studies the finite population mean-field control problems and their infinite population limit, and provide solutions of the ergodic control problems for some special cases.
@@
In the context of learning, [22, 23] study dynamic programming principle and Q learning methods directly for the infinite population control problem. The value functions and the Q functions are defined for the lifted problem, where the state process is the measure-valued mean-field flow. They consider dynamics without common noise, and thus the learning problem from the perspective of a coordinator becomes a deterministic one.
@@
[15] also considers the limit (infinite population) problem and studies different classes of policies that achieve optimal performance for the infinite population (limit problem) and focuses on Q learning methods for the problem after establishing the optimality of randomized feedback policies for the agents. The learning problem considers the state as the measure valued mean-field term and defines a learning problem over the set of probability measures where various approximations are considered to deal with the high dimension issues.
@@
[4, 3] have studied learning methods for the mean-field game and control problems from a joint lens. However, for the control setup, they consider a different control objective compared to the previously cited papers. In particular, they aim to optimize the asymptotic phase of the control problem where the agents are assumed to reach to their stationary distributions under joint symmetric policies. Furthermore, the agents only use their local state variables, and thus the objective is to find a stationary measure for the agents where the cost is minimized under this stationary regime. Since the agents only use their local state variables (and not the mean-field term) for their control, the authors can define a Q function over the finite state and action spaces of the agents.
@@
[34] consider a closely related problem to our setting, where they propose model-based learning methods for the mean-field control. Similar to [22, 23], they directly work with the infinite population dynamics without analyzing the approximation consistency between the finite-population dynamics and their infinite-population counterpart. Furthermore, they restrict the dynamics to the models with additive noise, and the optimality search is within deterministic and Lipschitz continuous controls.
@@
We also note that there are various studies that focus on the application of the mean-field modeling using numerical methods based on machine learning techniques, see e.g. the works by [37, 1, 29].
@@
In this paper, we will consider the learning problem using an alternative formulation where the state is represented as the measure valued mean-field term. To approximate this uncountable space, and the cost and transition functions, different from the previous works in the mean-field control setting, we will consider linear function approximation methods. These methods have been studied well for single agent discrete time stochastic control problems. We cite papers by [31, 16, 41, 27] in which reinforcement learning techniques are used to study Markov decision problems with continuous state spaces using linear function approximations.
@@
The dynamics for the model are presented as follows: suppose $N$ agents (decision-makers or controllers) act in a cooperative way to minimize a cost function, and the agents share a common state and an action space denoted by $\mathds{X}$ and $\mathds{U}$. We assume that $\mathds{X}$ and $\mathds{U}$ are finite. We refer the reader to the paper by [6], for finite approximations of mean-field control problems where the state and actions spaces of the agents are continuous. For any time step $t$, and agent $i\in\{1,\dots,N\}$ we have
@@
    the curse of coordination: even if the optimal team policy is found, its execution at the agent level requires coordination among the agents. In particular, the agents may need to follow asymmetric policies to achieve optimality, even though we assume full symmetry for the dynamics and the cost models. The following simple example from [9] shows that the agents may need to follow asymmetric policies to achieve optimality which requires coordination among the agents.
Example 1.1.
@@
A standard approach to deal with mean-field control problems when $N$ is large is to consider the infinite population problem, i.e. taking the limit $N\to\infty$. A propagation of chaos argument can be used to show that in the limit, the agents become asymptotically independent. Hence, the problem can be formulated from the perspective of a representative single agent. This approach is suitable to deal with coordination challenges, as the correlation between the agents vanish in the limit, and thus the symmetric policies can achieve optimal performance for the infinite population control problem. In particular, for Example 1.1 in the infinite population setting, the optimal policy is to follow a randomized policy such that $Pr(u=1)=Pr(u=0)=\frac{1}{2}$. We will introduce the limit problem in Section 1.5 and make the connections between the limit problem and the finite population problem rigorous.
@@
Recall that we assume that the state and action spaces of agents $\mathds{X,U}$ are finite (see [6] for finite approximations of continuous space mean-field control problems).
@@ 1.4 Measure Valued Formulation of the Finite Population Control Problem
For the remaining part of the paper, we will often consider an alternative formulation of the control problem for the finitely many agent case where the controlled process is the state distribution of the agents, rather than the state vector of the agents. We refer the reader to [6] for the full construction; in this section, we will give an overview.
@@
One can show that [6, 9] the empirical distributions of the states of agents $\mu_{t}$, and of the joint state and actions $\Theta_{t}$ define a controlled Markov chain such that
@@
Theorem 1 ([6]).
Under Assumption 1, for any ${\bf x}_{0}$ that satisfies $\mu_{{\bf x}_{0}}=\mu_{0}$, that is for any ${\bf x}_{0}$ with distribution $\mu_{0}$, we have that
@@ 1.5 Mean-field Limit Problem
The following result makes the connection between the finite population and the infinite population control problem rigorous [32, 5, 9].
Assumption 1.
@@
Theorem 2.
Under Assumption 1, the following holds:
@@ 1.6 Limitations of Full Decentralization
In words, the state distribution should be either be distributed equally over the state space $\{0,1\}$ or it should fully concentrate in state 00 for minimal stage-wise cost. One can check that the cost function satisfies Assumption 1 for some $K_{c}<\infty$ (e.g. $K_{c}=1$). For the dynamics we assume a deterministic model such that
@@ 2.1 Coordinated Learning with Linear Function Approximation for Infinitely Many Players
$P_{x}\subset{\mathcal{P}}(\mathds{X})$ denotes the set of probability measures which assign positive measure to a particular state $x\in\mathds{X}$ that are also in the training data for the mean-field terms. In particular, for a given $(x,u)$ pair, the kernel $\mathcal{T}(\cdot|x,u,\mu)$ and the cost $c(x,u,\mu)$ can be learned perfectly for every $\mu\in P_{x}$ with Assumption 3.
@@ 2.2 Independent Learning with Linear Function Approximation for Infinitely Many Players
Remark 3.
We can establish some sufficient conditions on the transition kernel of the system to test the ergodicity. We note that ([25, p 56, 3.1.1]) a sufficient condition is the following: there exists a mean-field state, say $\mu^{*}\in{\mathcal{P}}(\mathds{X})$ such that
@@
then one can set $\mu^{*}(x_{1})$ to be a value in this intersection independent of $\mu$. By doing this for all $x_{1}$, we can set a reachable $\mu^{*}$ from any $\mu\in{\mathcal{P}}(\mathds{X})$. As a result, if (20) holds for every $x_{1}$, then (19), and thus Assumption 4 can be shown to hold.
@@
We now define the trained measures, $P_{x}\in{\mathcal{P}}({\mathcal{P}}(\mathds{X}))$ for each $(x,u)$ pair based on the invariant measure for the mean-field flow. Under Assumption 3, that is assuming $Pr(u|x)=\int\gamma_{w}(u|x)P_{w}(dw)>0$ for all $(x,u)$ pairs, we can write
@@
Proof.
The proof can be found in Appendix A. ∎
Corollary 1.
Let Assumption 4 and Assumption 3 hold and let the learning rates be chosen such that $\alpha_{t}(x,u)=0$ unless $(X_{t},U_{t})=(x,u)$. Furthermore, $\sum_{t}\alpha_{t}(x,u)=\infty$ and $\sum_{t}\alpha_{t}^{2}(x,u)<\infty$ with probability one for all $(x,u)\in\mathds{X}\times\mathds{U}$. Then, the iterations given in (22) and (23) converge with probability 1. Furthermore, the limit points, say ${\bf\theta}^{*}(x,u)$ and ${\bf Q}^{*}_{(x,u)}(\cdot)$ are such that
@@
Proof.
We define the following stopping times
@@
For the iterations (22), Proposition 1 applies such that for each $(x,u)$, $v_{k}\equiv\theta_{\tau_{k}}(x,u)$, $k(\mu)\equiv{\bf\Phi}^{\intercal}(x,u,\mu)$ and $h(\mu)\equiv c(x,u,\mu)$ and finally the noise process $s_{k}\equiv\mu_{\tau_{k}}$. Note that ${\bf\Phi}$ and $c$ are assumed to be uniformly bounded which also agrees with the assumptions in Propositions 1. Furthermore, with the strong Markov property, $\mu_{\tau_{k}}$ is also a Markov chain which is sampled when the state-action pair is $(x,u)$. Thus, the invariant measure for the sampled process is $P_{x}$ as defined in (2.2).
For the iterations (23), Proposition 1 applies such that for each $(x,u)$ and each $j$, $v_{t}\equiv{\bf Q}^{j}_{\tau_{k}}(x,u)$, $k(\mu)\equiv{\bf\Phi}^{\intercal}(x,u,\mu)$ and $h(\mu)\equiv\mathds{1}_{\{X_{1}=x^{j}\}}$. We note that $X_{1}=f(x,u,\mu,w)$ (see (1)) where $w$ is the i.i.d. noise for the dynamics of agents. Thus, the noise process for iterations (23) can be taken to be the joint process $(\mu_{t},w_{t})$ where $\mu_{t}$ is an ergodic Markov process, and $w_{t}$ is an i.i.d. process. In particular, for every $(x,u)$ pair and for every $x^{j}$, if we consider the expectation over $(\mu,w)$ where $\mu\sim P_{x}(\cdot)$, we get
@@ 2.3 Learning for Finitely Many Players
and thus [25, p 56, 3.1.1] implies that the process ${\bf X}_{t}$ is geometrically ergodic.
The next result shows the convergence of the algorithm. Similar to the previous section, we first define the trained sets of mean-field terms for every $(x,u)$ pair using the stationary distribution of the mean-field terms. We assume that Assumption 3 holds for every policy $\gamma^{i}$ such that $\gamma^{i}(u|x)>0$ for every $(x,u)$ pair. Denoting the invariant distribution of the joint state process by $P(\bf x)$, for some $\mu^{N}\in{\mathcal{P}}_{N}(\mathds{X})$, for agent $i$, we can write
@@
Corollary 2.
Let Assumption 5 hold, and let Assumption 3 hold for each policy $\gamma^{i}$. Assume further that the learning rates of every agent satisfy the assumption of Corollary 1 such that $\sum_{t}\alpha_{t}(x,u)=\infty$ and $\sum_{t}\alpha_{t}^{2}(x,u)<\infty$ with probability one for all $(x,u)\in\mathds{X}\times\mathds{U}$. Then, the iterations given in (24) and (25) converge with probability one. Furthermore, for agent $i$, the limit points, say ${\bf\theta}^{*}_{(x,u)}$ and ${\bf Q}^{*}_{(x,u)}(\cdot)$ are such that
@@
Proof.
The proof is identical to the proof of Corollary 1, and is an application of Proposition 1. The only difference is the ergodicity of the mean-field process $\mu_{t}^{N}$, which does not require common randomness for exploration policies. ∎

### v3

@@ 1.1 Literature Review
Learning for multi agent control problems is a practically relevant and a challenging problem where there has been as a growing interest in recent years. A general solution methodology for multi-agent control problems is difficult to obtain and the solution, in general, is intractable except for special information structures between the agents. We refer the reader to the survey paper by Zhang et al. (2021) for a substantive summary of learning methods in the context of multi-agent decision making problems.
@@
In this paper, we study a particular case of multi-agent problems in which both the agents and their interactions are symmetric and homogeneous. For these mean-field type decision making problems, the agents are coupled only through the so-called mean-field term. These problems can be broadly divided into two categories; mean-field game problems where the agents are competitive and interested in optimizing their self objective functions, and mean-field control problems, where the agents are interested in a common objective function optimization. We cite some papers by Gomes and Saúde (2014); Carmona and Delarue (2013); Bensoussan et al. (2013); Tembine et al. (2013); Huang et al. (2007); Anahtarci et al. (2022); Elie et al. (2020); Fu et al. (2020); Guo et al. (2019); Perrin et al. (2020); Subramanian and Mahajan (2019); Saldi et al. (2018, 2019) and references therein, for papers in mean-field game setting. We do not discuss these in detail as our focus will be on mean-field control problems which are significantly different in both analysis and the nature of the problems of interest.
@@
For mean-field control problems, where the agents are cooperative and work together to minimize (or maximize) a common cost (or reward) function, see Bayraktar et al. (2018); Djete et al. (2022); Laurière and Pironneau (2014); Carmona and Laurière (2022); Pham and Wei (2017); Carmona and Laurière (2021); Germain et al. (2022); Bayraktar et al. (2023); Bayraktar and Zhang (2023) and references therein for the study of dynamic programming principle and learning methods in continuous time. In particular, we point out the papers Lacker (2017); Djete et al. (2022) which provide the justification for studying the centralized limit problem by rigorously connecting the large population decentralized setting and the infinite population limit problem.
@@
For papers studying mean-field control in discrete time, we refer the reader to Motte and Pham (2023); Bäuerle (2023); Gu et al. (2021, 2023); Motte and Pham (2022); Carmona et al. (2023). Motte and Pham (2023, 2022) study existence of solutions to the control problem in both infinite and finite population settings, and they rigorously establish the connection between the finite and infinite population problems. Bäuerle (2023) studies the finite population mean-field control problems and their infinite population limit, and provide solutions of the ergodic control problems for some special cases.
@@
In the context of learning, Gu et al. (2021, 2023) study dynamic programming principle and Q learning methods directly for the infinite population control problem. The value functions and the Q functions are defined for the lifted problem, where the state process is the measure-valued mean-field flow. They consider dynamics without common noise, and thus the learning problem from the perspective of a coordinator becomes a deterministic one.
@@
Carmona et al. (2023) also considers the limit (infinite population) problem and studies different classes of policies that achieve optimal performance for the infinite population (limit problem) and focuses on Q learning methods for the problem after establishing the optimality of randomized feedback policies for the agents. The learning problem considers the state as the measure valued mean-field term and defines a learning problem over the set of probability measures where various approximations are considered to deal with the high dimension issues.
@@
Angiuli et al. (2023, 2022) have studied learning methods for the mean-field game and control problems from a joint lens. However, for the control setup, they consider a different control objective compared to the previously cited papers. In particular, they aim to optimize the asymptotic phase of the control problem where the agents are assumed to reach to their stationary distributions under joint symmetric policies. Furthermore, the agents only use their local state variables, and thus the objective is to find a stationary measure for the agents where the cost is minimized under this stationary regime. Since the agents only use their local state variables (and not the mean-field term) for their control, the authors can define a Q function over the finite state and action spaces of the agents.
@@
Pásztor et al. (2023) consider a closely related problem to our setting, where they propose model-based learning methods for the mean-field control. Similar to Gu et al. (2021, 2023), they directly work with the infinite population dynamics without analyzing the approximation consistency between the finite-population dynamics and their infinite-population counterpart. Furthermore, they restrict the dynamics to the models with additive noise, and the optimality search is within deterministic and Lipschitz continuous controls.
@@
We also note that there are various studies that focus on the application of the mean-field modeling using numerical methods based on machine learning techniques, see e.g. the works by Ruthotto et al. (2020); Achdou et al. (2020); Lauriere et al. (2022).
@@
In this paper, we will consider the learning problem using an alternative formulation where the state is represented as the measure valued mean-field term. To approximate this uncountable space, and the cost and transition functions, different from the previous works in the mean-field control setting, we will consider linear function approximation methods. These methods have been studied well for single agent discrete time stochastic control problems. We cite papers by Melo et al. (2008); Carvalho et al. (2020); Szepesvári and Smart (2004); Jin et al. (2020); Meyn (2024) in which reinforcement learning techniques are used to study Markov decision problems with continuous state spaces using linear function approximations.
@@
The dynamics for the model are presented as follows: suppose $N$ agents (decision-makers or controllers) act in a cooperative way to minimize a cost function, and the agents share a common state and an action space denoted by $\mathds{X}$ and $\mathds{U}$. We assume that $\mathds{X}$ and $\mathds{U}$ are finite. We refer the reader to the paper by Bayraktar et al. (2025), for finite approximations of mean-field control problems where the state and actions spaces of the agents are continuous. For any time step $t$, and agent $i\in\{1,\dots,N\}$ we have
@@
    the curse of coordination: even if the optimal team policy is found, its execution at the agent level requires coordination among the agents. In particular, the agents may need to follow asymmetric policies to achieve optimality, even though we assume full symmetry for the dynamics and the cost models. The following simple example from Bayraktar and Kara (2024) shows that the agents may need to follow asymmetric policies to achieve optimality which requires coordination among the agents.
Example 1
@@
A standard approach to deal with mean-field control problems when $N$ is large is to consider the infinite population problem, i.e. taking the limit $N\to\infty$. A propagation of chaos argument can be used to show that in the limit, the agents become asymptotically independent. Hence, the problem can be formulated from the perspective of a representative single agent. This approach is suitable to deal with coordination challenges, as the correlation between the agents vanish in the limit, and thus the symmetric policies can achieve optimal performance for the infinite population control problem. In particular, for Example 1 in the infinite population setting, the optimal policy is to follow a randomized policy such that $Pr(u=1)=Pr(u=0)=\frac{1}{2}$. We will introduce the limit problem in Section 1.5 and make the connections between the limit problem and the finite population problem rigorous.
@@
Recall that we assume that the state and action spaces of agents $\mathds{X,U}$ are finite (see Bayraktar et al. (2025) for finite approximations of continuous space mean-field control problems).
@@ 1.4 Measure Valued Formulation of the Finite Population Control Problem
For the remaining part of the paper, we will often consider an alternative formulation of the control problem for the finitely many agent case where the controlled process is the state distribution of the agents, rather than the state vector of the agents. We refer the reader to Bayraktar et al. (2025) for the full construction; in this section, we will give an overview.
@@
One can show that Bayraktar et al. (2025); Bayraktar and Kara (2024) the empirical distributions of the states of agents $\mu_{t}$, and of the joint state and actions $\Theta_{t}$ define a controlled Markov chain such that
@@
Theorem 1 (Bayraktar et al. (2025))
Under Assumption 1.1, for any ${\bf x}_{0}$ that satisfies $\mu_{{\bf x}_{0}}=\mu_{0}$, that is for any ${\bf x}_{0}$ with distribution $\mu_{0}$, we have that
@@ 1.5 Mean-field Limit Problem
The following result makes the connection between the finite population and the infinite population control problem rigorous Motte and Pham (2022); Bäuerle (2023); Bayraktar and Kara (2024).
Assumption 1.1
@@
Theorem 2
Under Assumption 1.1, the following holds:
@@ 1.6 Limitations of Full Decentralization
In words, the state distribution should be either be distributed equally over the state space $\{0,1\}$ or it should fully concentrate in state 00 for minimal stage-wise cost. One can check that the cost function satisfies Assumption 1.1 for some $K_{c}<\infty$ (e.g. $K_{c}=1$). For the dynamics we assume a deterministic model such that
@@ 2.1 Coordinated Learning with Linear Function Approximation for Infinitely Many Players
$P_{x}\subset{\mathcal{P}}(\mathds{X})$ denotes the set of probability measures which assign positive measure to a particular state $x\in\mathds{X}$ that are also in the training data for the mean-field terms. In particular, for a given $(x,u)$ pair, the kernel $\mathcal{T}(\cdot|x,u,\mu)$ and the cost $c(x,u,\mu)$ can be learned perfectly for every $\mu\in P_{x}$ with Assumption 2.2.
@@ 2.2 Independent Learning with Linear Function Approximation for Infinitely Many Players
Remark 5
We can establish some sufficient conditions on the transition kernel of the system to test the ergodicity. We note that ((Hernandez-Lerma, 2012, p 56, 3.1.1)) a sufficient condition is the following: there exists a mean-field state, say $\mu^{*}\in{\mathcal{P}}(\mathds{X})$ such that
@@
then one can set $\mu^{*}(x_{1})$ to be a value in this intersection independent of $\mu$. By doing this for all $x_{1}$, we can set a reachable $\mu^{*}$ from any $\mu\in{\mathcal{P}}(\mathds{X})$. As a result, if (20) holds for every $x_{1}$, then (19), and thus Assumption 2.3 can be shown to hold.
@@
We now define the trained measures, $P_{x}\in{\mathcal{P}}({\mathcal{P}}(\mathds{X}))$ for each $(x,u)$ pair based on the invariant measure for the mean-field flow. Under Assumption 2.2, that is assuming $Pr(u|x)=\int\gamma_{w}(u|x)P_{w}(dw)>0$ for all $(x,u)$ pairs, we can write
@@
Proof  The proof can be found in Appendix A.
Corollary 6
Let Assumption 2.3 and Assumption 2.2 hold and let the learning rates be chosen such that $\alpha_{t}(x,u)=0$ unless $(X_{t},U_{t})=(x,u)$. Furthermore, $\sum_{t}\alpha_{t}(x,u)=\infty$ and $\sum_{t}\alpha_{t}^{2}(x,u)<\infty$ with probability one for all $(x,u)\in\mathds{X}\times\mathds{U}$. Then, the iterations given in (22) and (23) converge with probability 1. Furthermore, the limit points, say ${\bf\theta}^{*}(x,u)$ and ${\bf Q}^{*}_{(x,u)}(\cdot)$ are such that
@@
Proof  We define the following stopping times
@@
For the iterations (22), Proposition 2.1 applies such that for each $(x,u)$, $v_{k}\equiv\theta_{\tau_{k}}(x,u)$, $k(\mu)\equiv{\bf\Phi}^{\intercal}(x,u,\mu)$ and $h(\mu)\equiv c(x,u,\mu)$ and finally the noise process $s_{k}\equiv\mu_{\tau_{k}}$. Note that ${\bf\Phi}$ and $c$ are assumed to be uniformly bounded which also agrees with the assumptions in Propositions 2.1. Furthermore, with the strong Markov property, $\mu_{\tau_{k}}$ is also a Markov chain which is sampled when the state-action pair is $(x,u)$. Thus, the invariant measure for the sampled process is $P_{x}$ as defined in (2.2).
For the iterations (23), Proposition 2.1 applies such that for each $(x,u)$ and each $j$, $v_{t}\equiv{\bf Q}^{j}_{\tau_{k}}(x,u)$, $k(\mu)\equiv{\bf\Phi}^{\intercal}(x,u,\mu)$ and $h(\mu)\equiv\mathds{1}_{\{X_{1}=x^{j}\}}$. We note that $X_{1}=f(x,u,\mu,w)$ (see (1)) where $w$ is the i.i.d. noise for the dynamics of agents. Thus, the noise process for iterations (23) can be taken to be the joint process $(\mu_{t},w_{t})$ where $\mu_{t}$ is an ergodic Markov process, and $w_{t}$ is an i.i.d. process. In particular, for every $(x,u)$ pair and for every $x^{j}$, if we consider the expectation over $(\mu,w)$ where $\mu\sim P_{x}(\cdot)$, we get
@@ 2.3 Learning for Finitely Many Players
and thus (Hernandez-Lerma, 2012, p 56, 3.1.1) implies that the process ${\bf X}_{t}$ is geometrically ergodic.
The next result shows the convergence of the algorithm. Similar to the previous section, we first define the trained sets of mean-field terms for every $(x,u)$ pair using the stationary distribution of the mean-field terms. We assume that Assumption 2.2 holds for every policy $\gamma^{i}$ such that $\gamma^{i}(u|x)>0$ for every $(x,u)$ pair. Denoting the invariant distribution of the joint state process by $P(\bf x)$, for some $\mu^{N}\in{\mathcal{P}}_{N}(\mathds{X})$, for agent $i$, we can write
@@
Corollary 9
Let Assumption 2.4 hold, and let Assumption 2.2 hold for each policy $\gamma^{i}$. Assume further that the learning rates of every agent satisfy the assumption of Corollary 6 such that $\sum_{t}\alpha_{t}(x,u)=\infty$ and $\sum_{t}\alpha_{t}^{2}(x,u)<\infty$ with probability one for all $(x,u)\in\mathds{X}\times\mathds{U}$. Then, the iterations given in (24) and (25) converge with probability one. Furthermore, for agent $i$, the limit points, say ${\bf\theta}^{*}_{(x,u)}$ and ${\bf Q}^{*}_{(x,u)}(\cdot)$ are such that
@@
Proof  The proof is identical to the proof of Corollary 6, and is an application of Proposition 2.1. The only difference is the ergodicity of the mean-field process $\mu_{t}^{N}$, which does not require common randomness for exploration policies.